OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 2x + y = 5 3x – 2y = 25

OpenStudy (nubeer):

y = 5-2x call it equation 1 now substitute this in the other equation 3x - 2(5-2x) = 25 3x -10 +4x = 25 7x =35 x = 5 now substitute this x in equation 1 y = 5-2(5) y = -5